Suppose v1,v2,........,vm is a linearly independent in V and w∈V. Show that v1,v2,........,vm is linearly independent if and only if w∉span(v1,v2,........,vm )

1
Expert's answer
2021-05-11T13:58:45-0400

Suppose that "w\in \text{Span}(v_1,...,v_m)", then we have "w=\sum \alpha_i v_i" and thus "\sum \alpha_iv_i-w=0" is a non-trivial (as at least the coefficient before "w" is "-1\neq 0" ) linear combination of "(v_1,...,v_m,w)" that gives zero and thus "(v_1,...,v_m,w)" is not linearly independent.

Now let's suppose that "(v_1,...,v_m,w)" is linearly dependent, then there exists a non-trivial linear combination such that "\sum \alpha_i v_i +bw=0". We can not have "b=0" as in this case we would have "\sum \alpha_i v_i=0" for sum non-trivial combination "(\alpha_i)". Therefore we have "w=\sum(-\alpha_i/b)v_i \in \text{Span}(v_1,...,v_m)".

"\neg((v_1,...,v_m,w) \text{ are linearly independent}) \leftrightarrow \neg (w\notin \text{Span}(v_1,...,v_m))" and therefore we have (by contraposition)

"((v_1,...,v_m,w) \text{ are linearly independent}) \leftrightarrow (w\notin \text{Span}(v_1,...,v_m))"